OpenStudy (anonymous):

y = 2/x^2 +3 Find f(x) and g(x) so the function can be expressed as y = f(g(x))

OpenStudy (anonymous):

difrentiate it ...

OpenStudy (anonymous):

I'd say the "inside" or nested function is x^2 and the outside function is 2/x + 3 That makes f(x) = 2/x + 3 and g(x0=x^2 Does that make sense?